OpenStudy (anonymous):

Equation of a line tangent to the curve xy = √(xy – x) + 1 at the point (1, 2)? I have some work done...

OpenStudy (anonymous):

So I did manage to differentiate it: (x(3-2xy))/(1-3y+2xy^2), and then I plugged in my values for x and y, (1,2) and got -1/3. Am I on track so far? What would the equation of the tangent lint be then?

OpenStudy (anonymous):

*line

OpenStudy (dumbcow):

well if you have the slope (m)...then equation of tangent line is: y - y1 = m(x -x1) x1,y1 is point line goes through (1,2) y-2 = -1/3(x-1)

OpenStudy (anonymous):

That's my problem at this point. When I isolate y, I get (7-x)/3, which is not equivalent to any of my answer choices.

OpenStudy (anonymous):

what did you get when you differentiate ?

OpenStudy (anonymous):

(x(3-2xy))/(1-3y+2xy^2) Did I do something wrong differentiating? (I'm not that great with implicit stuff, so that might be the case.)

OpenStudy (dumbcow):

yes, sorry i didn't check before....your derivative is wrong when you have to do implicit differentiation, don't forget to use product rule \[(xy)' = xy' + x'y = x \frac{dy}{dx} + y\]

OpenStudy (dumbcow):

\[\rightarrow x \frac{dy}{dx} +y = \frac{x \frac{dy}{dx} +y -1}{2 \sqrt{xy-x}}\]

OpenStudy (dumbcow):

\[\rightarrow \frac{dy}{dx} = \frac{-2y \sqrt{xy-x} +y-1}{x(2 \sqrt{xy-x} -1)}\]

OpenStudy (dumbcow):

were you able to follow my work? got the answer now

OpenStudy (anonymous):

Yes I think so -- plug and chug: (-2(2)sqrt(1(2)-1)+(2)-1)/(1(2sqrt(1(2)-1)-1)) = 3 is what I got. Plugging that in and solving for y gives me 3x-1, but that's still not in the answer choices... It's close to option A. I might just go with that.

OpenStudy (dumbcow):

haha yes that would be a good educated guess. that is the answer...slope should come out to be -3 when you plug in (1,2)

OpenStudy (anonymous):

Oh durr, it's -3, and that changes everything It's A, thank you for putting me on track with the differentiation!

OpenStudy (dumbcow):

no problem